The temperatures on the inside and outside surface of a wall are 20 degrees * C and 5 degrees * C respectively The wall is 8 m wide4 m high and 0.2 m thick. The thermal conductivity is 0.3W / m * o * K


what is the heat flux through the wall?


1
Expert's answer
2021-01-21T03:50:52-0500

k= 0.3 W/m K

ΔT = 20 – 5 = 15 ºC

"S = 8 \times 4 = 32 \;m^2"

A = 0.2 m

"q = \frac{ΔT}{A/kS} \\

= \frac{15}{0.2/(0.3 \times 32)} \\

= 720 \;W"

Answer: 720 W